OpenStudy (anonymous):

solve -10over13 = -2 1over2 b will give medal to best answer.

OpenStudy (anonymous):

1. b = -4/13 2. b=1 12/13 3. b= 4/13 4. b = -1 12/13

OpenStudy (anonymous):

Can you specify more on what it looks like? I cannot tell what is on the right side. -2*0.5b?

OpenStudy (anonymous):

\[\frac{ -10 }{ 13 } = -2 \frac{ 1 }{ 2 } b\]

OpenStudy (anonymous):

@Potatocracy

OpenStudy (anonymous):

Thank you. To get the right side condensed you can pretend -2 is -2/1 and multiply it to 1/2. This will get you -1b. Then you divide out the negative and it will equal 10/13 but that is not an answer. Maybe there is a different operation between -2 and 1/2?

OpenStudy (anonymous):

ok, so would the answer be 1? b = -4/13 @Potatocracy

OpenStudy (anonymous):

From the equation you wrote the answer is 10/13, so I am pretty sure you wrote it wrong or something.

OpenStudy (anonymous):

lol, nope. tell that to my teacher.

OpenStudy (anonymous):

Lol. Well sorry then I can't help you :3